### v2

# MAVIL: Design of a Multidimensional Assessment of Visual Data Literacy and its Application in a Representative Survey

Abstract
The ability to read, interpret, and critique data visualizations has mainly been assessed using data visualization tasks like value retrieval. Although evidence on different facets of Visual Data Literacy (VDL) is well established in visualization research and includes numeracy, graph familiarity, or aesthetic elements, they have not been sufficiently considered in ability assessments. Here, VDL is approached as a multidimensional ability whose facets can be partially assessed with self-perceived ability ratings. We introduce an assessment in which VDL is deconstructed as a process of understanding, in reference to frameworks from the learning sciences. MAVIL (Multidimensional Assessment of Visual Data Literacy) comprises six ability dimensions: Comprehending, Decoding, Aestheticizing, Critiquing, Reading, and Contextualizing. MAVIL was designed for general audiences, with validation supported by a survey (N=438) representative of a German-speaking country’s age groups (18-74 years) and male-female gender split. The survey mirrors the population’s VDL—encompassing self-assessment, performance measures and open-ended visualization critique—and shows the understanding of two climate data visualizations, a line and bar chart. For instance, we found that about a quarter have self-assessed deficits in comprehending simple data units, roughly one in five people feel unfamiliar with each displayed chart type and that self-assessed numeracy relates significantly to data reading performance (p=$0.0004$).
keywords:
Literacy, Data and Knowledge Visualization, Data Communications Aspects, Test Design, Self-Assessment
1 Motivation
Visual Data Literacy (VDL) is the human ability to read, interpret, and critique visual data representations [78]. Even though insights into different dimensions of this ability are well established in visualization research and include numeracy [69], graph familiarity [31], or aesthetic elements [70], they remain underutilized in ability assessments. We introduce MAVIL (Multidimensional Assessment of VDL), an approach that integrates self-perceived ability ratings and provides different and more nuanced insights into the understanding of tested visualizations. Through qualitative pilots (N=$8$) and survey validation (N=$438$), we demonstrate its potential for providing insights into VDL facets.
A common assessment strategy for VDL relies on data visualization tasks like value retrieval. Exemplary of such assessment structures are the Visualization Literacy Assessment Test (VLAT) [78], its briefer version, the Mini-VLAT [67], or the Critical Thinking Assessment for Literacy in Visualizations (CALVI) [34]. MAVIL complements existing approaches by incorporating self-perceived abilities, enabling an assessment of VDL that considers both task-based performance and how visualization readers perceive their own skills. By incorporating self-assessment ratings, MAVIL captures nuances that task-based tests may miss, such as how trustworthy a visualization seems or self-indicated difficulties linked to specific graph elements. MAVIL is composed of six ability dimensions:
  • Comprehending: Overall Impression and Abstract Thinking,
  • Decoding: Graph Component Understanding,
  • Aestheticizing: Aesthetic Perception,
  • Critiquing: Visualization Criticism,
  • Reading: Low-level Data Reading Tasks,
  • Contextualizing: Numeracy and Topic Knowledge.
Designed for general audiences, MAVIL addresses some limitations of prior assessments that strongly focused on expert audiences [31, 40] or specific groups, such as students [13, 30]. Given the increasing need for data literacy—and not everyone being equally data-savvy [52]—our assessment is designed to engage a broad audience, including people of varying ages and without expertise in data visualization.
Here, we showcase the design of MAVIL and the analysis of its implementation in a survey with representative scope for a German-speaking country regarding age groups (18-74 years) and the male-female gender split within them. The survey evaluation illustrates how MAVIL can reveal specific ability dimensions, such as perceived deficits, of visualization readers in a given context. Moreover, it highlights how respondents perceived two tested visualizations, a line chart and bar chart on climate data, chosen for their simplicity, familiarity, and importance in conveying critical information [60, 66, 73].
We found that $48\%$ of respondents (ages 18-74) in the surveyed country make mistakes with the simple line and bar chart. About a quarter of respondents found simple data units, e.g., millimeters, in the surveyed charts incomprehensible. Further, about $12\%$ felt unfamiliar with both displayed graph types. Concerning aesthetic perception, there was a tendency to find the visualization pretty or well-designed and simultaneously rate it as trustworthy. The assessment also included optional text fields for visualization criticism, in which $37\%$ of survey respondents reported at least one point of critique. The qualitative evaluation of text inputs revealed not only respondents’ critical abilities but also which visualization elements can irritate. In our survey, respondents’ self-assessed numeracy level relates significantly (p=$0.004$) to data reading performance and all other dimensions of VDL proposed in this paper.
Ability Dimension
in MAVIL
Perspective
Ability Examples Related to
Visual Data Literacy
Related Facet of
Understanding [80]
Exemplary Findings From
Our Representative Survey
Dim1: Comprehending Top-Down
Summarize visualized content,
recall information [7, 40]
Explanation
$76\%$ feel confident in
summarizing
Dim2: Decoding Bottom-Up
Examine color usage, analyze
graph elements [70, 31]
Interpretation
$12\%$ feel unfamiliar with both
line and bar charts
Dim3: Aestheticizing Bottom-Up
Decide upon trustworthiness,
judge visual appeal [18, 41]
Empathy
Perceived beauty was linked
to trust
Dim4: Critiquing Bottom-Up
Articulate criticism, make
design suggestions [10, 17]
Perspective
$37\%$ submitted meaningful
critique in optional inputs
Dim5: Reading Top-Down
Solve data reading tasks
[78, 67]
Interpretation
$48\%$ made errors in data
reading tasks
Dim6: Contextualizing Top-Down
Utilize statistical skill, put
visualization in context [81, 9]
Application
Self-assessed numeracy
correlated with data reading
Table 1: Visual Data Literacy dimensions in MAVIL, developed from either a viewer-focused (top-down) or visualization-focused (bottom-up) perspective. Each dimension is illustrated by example activities and aligned with a corresponding Facet of Understanding [80]. Additionally, we share insights from our representative survey where we tested MAVIL, which is evaluated in detail in Section 5.
2 Background
Visual Data Literacy (VDL) has been described as the ability to read, interpret, and critically deal with data visualizations [78]. We developed MAVIL’s ability dimensions through a theory-driven approach, drawing on relevant studies and definitions, and combining a bipartite framework with learning sciences concepts that define understanding as a multifaceted process. Creating an assessment that accounts for the multi-dimensional composition of VDL was deemed necessary because current assessments are dominantly task-based, and therefore, they are perspectively limited in their ability display. We highlight studies using self-perceived abilities in relation to performance, arguing that such ratings can offer unique insights into VDL assessments beyond established methods.
2.1 Visual Data Literacy as a Multidimensional Ability
Visual Literacy has been explored as an intellectual and social phenomenon since the 1960s, with early contributions emerging in fields such as education, communication, visual arts, and psychology [58], addressing both audience-related factors and properties of visual design. In MAVIL, we adopt a bipartite framework, that covers top-down (representing the visualization reader) and bottom-up (representing the visualization) aspects of Visual Literacy. Additionally, we draw on insights from the learning sciences, which break complex processes, such as understanding, down into sub-skills.
2.1.1 Considering Viewer and Visualization Perspectives
Visual Literacy can be viewed through a bipartite perspective, distinguishing between top-down and bottom-up aspects [42]. Top-down characteristics take the perspective of visualization readers and are shaped by factors such as prior beliefs, demographic characteristics, and personal skills. Here, numeracy, for instance, has been identified as a critical skill influencing data reading accuracy and knowledge gains [81]. In contrast, bottom-up characteristics focus on the visualization itself, including its content or aesthetic structure [70]. We argue that when assessing Visual Data Literacy, it is valuable to include dimensions that capture the influence of visualization design on audience responses. Visualization techniques, such as color choices or the depiction of uncertainty, have been shown to directly influence how audiences retrieve information and take decisions [61, 79, 25].
Therefore, MAVIL incorporates both top-down and bottom-up aspects by aligning ability dimensions with either a viewer or visualization focus, as shown in Table 1. For example, Contextualizing (Dimension 6) represents top-down characteristics, and Decoding (Dimension 2) relates to bottom-up characteristics. For a full description of dimensions refer to Table 1. We further expanded this two-level approach by integrating frameworks that treat understanding as a multifaceted construct.
2.1.2 Deconstructing Visual Data Literacy into Ability Facets
Insights from learning sciences, such as the framework The Six Facets of Understanding [80], break the process of understanding down into different ability categories: explaining, interpreting, applying, empathizing, having perspective, and self-knowledge. Accordingly, a respondent who can summarize the content of a data visualization (MAVIL Q1c, for an overview of questions, see Figure 1) would be able to explain. If a respondent solves a data reading task (Q5a-Q5d), this is an act of interpretation. Deciding upon the trustworthiness of a visualization (Q3d) requires empathizing with its creator’s intent. Someone who can contextualize displayed data (Q6b) can apply. Critiquing or suggesting design changes (Q4b/Q4c) demonstrates perspective-taking. Self-assessed numeracy (Q6b) or interest in climate change (Q6a) means applying existing skills or knowledge.
Similarly, Bloom’s Taxonomy of Educational Objectives for Knowledge-Based Goals [4] deconstructs knowledge and learning into ability levels across cognitive, affective, and psychomotor domains, which are most commonly used for cognitive intents [53]. According to this framework, memorizing is followed by understanding, application, then analysis, evaluation, and creation. Applied to Visual Data Literacy, a respondent who recalls the information of a visualization (Q1c) remembers it. Someone who integrates information into their prior knowledge (Q1b) can understand. Solving a data reading task (Q5a-Q5d) demonstrates application. Examining color encoding (Q2b) involves analysis. Justifying a point of critique (Q4d/Q4e), represents evaluation. A respondent who suggests a design change (Q4b/Q4c) would be able to create and develop a visualization further. Bloom’s Taxonomy has, for example, been adapted to assess literacy with parallel coordinates plots [68].
Following these approaches from the learning sciences, we built MAVIL dimensions by linking facets of understanding to well-established Visual Data Literacy ability examples, as exemplified in Table 1.
2.2 Limitations of Prior Literacy Assessments
Prior literacy assessments of data visualizations often rely on task-based approaches, focus on a single data visualization type, or target specific audiences. Examples include treemap and parallel coordinate plot literacy tests [30, 27, 28]. Several studies addressed educational contexts, such as qualitative principles for assessing visual literacy in schools [13], while others examine expert understanding of climate data visualizations [31]. Despite these efforts, the utilized measurement approaches remain largely one-dimensional. Moreover, as noted in prior critiques, general audiences with varying ability levels are still underrepresented in visualization research [71, 6, 5].
In 2017, Lee et al. [78] introduced a Visualization Literacy Assessment Test (VLAT), a $53$-item multiple-choice test covering $12$ data visualizations. It is specialized in measuring performance for simple visualization tasks like value retrieval and has been implemented or adapted multiple times [62, 43]. While effective for measuring task-based performance, VLAT’s reliance on correct/incorrect responses limits its ability to capture broader dimensions of Visual Data Literacy, including insights into where deficits may lay and what visualization readers perceive as challenging. VLAT was recently condensed into the Mini-VLAT [67] to improve accessibility by reducing its original 2020+ minute duration. While this may improve accessibility, it still emphasizes numerical accuracy over other skills. The Critical Thinking Assessment for Literacy in Visualizations (CALVI) [34] extends VLAT by shifting its assessment focus on critical ability as part of Visualization Literacy. Comprising 45 items across nine chart types with intentional design flaws, CALVI meaningfully expands the concept of ability but remains focused on task performance. Other assessments ask respondents to construct visualizations themselves, such as in the AVEC assessment [35] or studies on parallel coordinate plot literacy [68], yet all of these remain mainly performance-based.
2.3 Self-Perception in Ability Assessments
Recently, PREVis, a scale for perceived readability in data visualizations, was introduced by Cabouat et al.[12] and exemplifies how self-assessment may provide insights into literacy. Using Likert scales, it measures four facets of perceived readability: understanding, visualization layout, data value readability, and perception of data patterns. PREVis is built on a strong theoretical foundation and provides a valuable entry point for respondents to self-assess the aspect of perceived readability [12], which we consider part of their visualization understanding. MAVIL overlaps with some PREVis aspects, such as perceived understanding (see MAVIL Q1a in Figure 1) and identifying visual elements (Q2a), but provides additional insights. For instance, our survey, detailed in Section 5, found that $22\%$ of respondents struggled to comprehend "millimeters", and $28\%$ found "millions of tons" unclear (Q2a). PREVis would not capture this hurdle to comprehension.
Self-assessment assumes that human traits and skills are quantifiable and measurable [77] and can complement task-based tests to support more nuanced approaches to measuring abilities [1, 32]. In our study, self-assessment captures how visualization readers perceive facets of their Visual Data Literacy. We hypothesize that these self-perception ratings can reveal nuanced insights into the ability and may correlate with performance in data reading tasks. Similar approaches have been widely applied in psychology, education, and health research: self-concepts of talent have been compared to performance [3], self-perceptions have been linked to academic achievement [76, 36], perceived coping abilities have been shown to predict smoking cessation outcomes [63]. Further, self-perception is recognized as a survey approach in political science, i.e., to gain insights into respondents’ stances and beliefs [59, 50].
While self-perceived ability ratings represent a different measurement type than task-based performance metrics and are not as "objective" or precisely quantifiable, they can still be predictive [2]. Rather than replacing established assessment tools like VLAT [78], we aim to integrate self-perceived ability ratings to enhance Visual Data Literacy assessments. Beyond measuring what people can do, self-assessment sheds light on how respondents perceive their interactions with visualizations. In literacy assessments, factors related to the latter, such as how personal beliefs, trust, or emotional responses influence information retrieval [21, 33], remain underexplored.
3 MAVIL: A Multidimensional VDL Assessment
MAVIL takes a multidimensional approach to Visual Data Literacy assessment, ensuring that all developed questions, including ones on self-perceived ability, are grounded in theoretical frameworks and empirical validation. Four out of six ability dimensions (Dim1-Dim3, Dim6) rely on self-perceived ability ratings, such as Comprehending with Q1a: "I think I understood the visualization". These ratings use five-point Likert scales, indicating levels of Visual Literacy or likelihood of correct interpretation. The according questions are based on a selection of literacy theories and current studies on data visualization understanding. One dimension (Dim5) covers low-level data visualization tasks, which relate to the tested visualization examples, and there are also optional text inputs (Dim4) for visualization critique. An overview of MAVIL’s ability dimensions and question formats is provided in Figure 1, with the dimensions outlined below in the order they were assessed, without implying a hierarchy.
Dimension 1: Comprehending. First, self-perception of overall graph comprehension, including abstract thinking, is surveyed. The general impressions of the provided data visualization example are addressed (Q1a/Q1b). Further, following a definition of deductive and inductive abstract thinking as an indicator of a high ability level [7], respondents are asked if they could easily summarize the content of the visualization to someone (Q1c).
Dimension 2: Decoding. Dimension 2 focuses on self-perceived decoding of graph specifics, such as graph elements and familiarity with chart types. The comprehensibility of color usage (Q2b), like the comprehensibility of depicted measurement units (Q2a), is queried. This reflects the understanding that visualization comprehension depends on both structural/content-related features and visual/aesthetic design elements [70]. Moreover, graph familiarity is said to be beneficial for reading data correctly [31], and therefore also investigated (Q2c).
Dimension 3: Aestheticizing. Dimension 3 addresses self-assessed aesthetic perception in relation to Visual Data Literacy. Here, the BeauVis scale [41] was adapted (Q3a-Q3c). This scale consists of singular, English-language adjectives, all synonymous and intended to iterate on aesthetic perception. We translated BeauVis’ descriptive terms into German, the primary language of the surveyed country, which was done in a language-sensitive manner to convey their initial idea. The questions on aesthetic perception are supplemented with a question on perceived trustworthiness (Q3d), as attributing beauty to a visualization may favorably bias its interpretation [18].
Figure 1: Structure of the MAVIL assessment. Questions are grouped into six Visual Data Literacy dimensions (Dim1–Dim6), alongside demographic questions. Question formats include Likert scale ratings, text fields, true/false tasks, and multiple-choice items. For the full question catalog, see the supplementary material.
(a) Data Visualization 1: A line chart showing sea level change over time with associated data reading tasks (Q5a, Q5b) and marked correct answers.
(b) Data Visualization 2: A bar chart comparing $\text{CO}_{2}$ emissions across selected countries (1990–2018) with tasks (Q5c, Q5d) and indicated correct answers.
Figure 2: Data visualizations used in the MAVIL survey with a German-speaking audience. Figure (a) shows an official English version of the tested line chart, while Figure (b) displays the orginial bar chart used in the survey, which is only available in German. See Section 4.1.1 for details.
Dimension 4: Critiquing. Dimension 4 centers on critical ability and understanding of visualization construction. Respondents are asked to articulate specific feedback on the data visualization. This includes indicating whether they have any erroneous findings to report (Q4a) or specific amendment proposals to share (Q4b/Q4c), as literacy covers the ability to question an object and to seek out possible changes [17]. This also plays into the idea of Visual Literacy as a grouping of acquired abilities and skills covering not only reading but also creating visualizations [10]. Respondents were offered subsequent optional text fields asking for visualization strengths and weaknesses (Q4d/Q4e).
Dimension 5: Reading. Dimension 5 complements MAVIL with task-based assessment questions by presenting data reading exercises (Q5a-Q5d), as this type of assessment question is a well established performance measure for literacy (see Section 2.2). Similar to related styles of assessments, such as in VLAT [78], these tasks directly refer to the data content of the surveyed visualization, as exemplified through true-or-false questions implemented for our surveyed visualizations in Figure 2.
Dimension 6: Contextualizing. Finally, two contextual ability facets closely related to Visual Data Literacy are surveyed: self-perception of numeracy skills and topic knowledge. A higher level of numeracy would indicate a higher likeliness of reading information from a data visualization correctly [69, 81], which is why we asked respondents rate their own numeracy level (Q6b). Beyond addressing this particular skill, personal interest in the displayed content is questioned (Q6a), as it may influence the understanding of visually represented data [31].
Even though most questions are generally phrased, some reference the data visualization example directly to exemplify parts of the questions to respondents. This is the case for data reading tasks (Q5a-Q5d) bound to the visualization content. Specifications were also made to Q2a by listing measurement units and Q2c by naming the chart type. In Q6a, respondents were asked to rate personal interest in climate-related topics.
4 Methodology
The design, development, and testing of MAVIL was split into four phases. At first, questions were developed in a theory-driven approach (for detail see Section 3) and placed within a digital survey format. Then, in two phases of qualitative testing, the assessment was presented to eight participants. Semi-structured interviews allowed us to gain insights into the validity of MAVIL questions, which were followed by iterative development. Finally, the assessment was completed by 438 respondents representing a German-speaking country’s age groups, from 18 to 74 years, and male-female gender distribution. MAVIL was developed and tested in the country’s primary language. The English version presented here is a translation and has not been separately tested. The original questions are made available at an anonymized link (to be provided in the camera-ready version).
4.1 MAVIL in Practice: Designing a Survey
MAVIL was implemented using SoSci Survey, a survey design and implementation software free for academic purposes [54]. The assessment was tested with four data visualizations throughout the pilots and two of them were used in the survey conduction, with each respondent always encountering two visualizations to minimize bias from any single visualization. To prevent order effects [8], automated urn-drawing method randomized the display sequence. This section discusses the rationale behind our visualization selection for testing and outlines survey tasks, which are fully detailed in the supplementary question catalog.
4.1.1 Selection of Visualization Examples on Climate Data
Data visualizations play a crucial role in scientific communication, such as when conveying critical information on the climate discourse [60, 66, 72]. While they can be effective tools for conveying complex information [65, 31], even the Intergovernmental Panel on Climate Change has acknowledged challenges in making them comprehensible to broader audiences [56, 40]. Studying their effectiveness is essential to understanding audience engagement with climate data better and ensuring trust in climate science [73].
The selection of data visualizations for MAVIL testing followed a number of criteria: they had to present climate-related data to match our topic setting, be simple enough to encourage engagement, they had to differ in chart type and remain static rather than interactive. Additionally, all visualizations were licensed under Creative Commons (CC BY-ND) for public use and credited accordingly. Their suitability for a general audience was judged during short feedback exchanges with visualization researchers.
Based on these criteria, we selected two visualizations with climate data for our survey, in particular a line chart and a bar chart (see Figure 2) were selected as they are among the most widely used visualization types [61], ensuring that respondents could engage with them regardless of prior expertise. The line chart (2(a)) shows sea level rise from 1993 to 2021, with an annotation highlighting an average annual increase of $3.5$ millimeters. The line is shaded with a light blue area labeled "Divergence," emphasizing the rate of increase. The bar chart (2(b)) compares $\text{CO}_{2}$ emissions of $10$ countries in 1990 and 2018, using color-coded bars and percentage changes. Each country’s flag is displayed and a $\text{CO}_{2}$ cloud icon is placed in the middle of the chart. During testing, both charts were in German, as this was the primary language of the surveyed audience.
4.1.2 Survey Tasks
Across all testing phases, respondents were presented with two data visualizations, meaning that Dimensions 1, 2, 3, and 4 were applied once per visualization example. Questions on Contextualizing appeared only once. Dimension 5, which addresses Reading, featured four tasks in true-or-false mode, two per visualization. Correct data reading responses across visualization examples were added to a Data Reading Score (DRS) for the analysis. An additional data reading task was designed by us and posed in reference to the line chart as an attention check to ensure survey data quality. This question had to be answered in an obligatory text field and functioned as an attention test.
To cater to a general audience, there are no prerequisites for taking the survey. Respondents are first informed about the study’s purpose of investigating the understanding of climate data visualizations. They have to confirm their consent to participate in the study and that they were not underage. We aimed to keep instructions throughout the survey concise but informative. Participants are asked to engage with two data visualizations, focusing on both content and overall impression. After viewing each visualization, they proceed to a set of related questions. Whenever a question refers directly to a visualization, the corresponding chart is displayed beneath it. Self-perceived ability ratings are presented on a five-point scale ($1$ = I disagree, $2$ = I somewhat disagree, $3$ = Undecided, $4$ = I somewhat agree, $5$ = I agree). For analysis, ratings of $1$ and $2$ were grouped to represent disagreement, and $4$ and $5$ as agreement. A rating of $3$ was considered a neutral answer. In Dimension 4 on Critiquing, respondents could provide optional text input, which was evaluated qualitatively by grouping responses into categories.
4.2 Qualitative Testing: Survey Pilots
Qualitative survey testing was carried out with eight respondents who filled out the assessment in the presence of a researcher. In semi-structured interviews, respondents described their interpretation of the two data visualization examples and discussed their experience answering MAVIL questions.
4.2.1 Purpose and Procedure of Survey Pilots
The survey pilots supported sharpening and validating MAVIL questions. By analyzing the respondents’ assessment experience, including question understanding, the feasibility of the questions could be qualitatively evaluated and then iterated upon. The insights were particularly productive for fine-tuning question phrasing. Respondents were recruited purposively through professional and personal relations, aiming for diverse educational and professional backgrounds. Six out of eight respondents had no expertise in climate data visualizations, one respondent was familiar with data visualizations due to their field of work, and one was deliberately a visualization researcher. The survey pilots were split into two phases, each consisting of four pilots. Sessions lasted between 28 and 47 minutes (median: 35 min). Interviews were audio-recorded with the consent of all participants and then transcribed.
4.2.2 Analysis of Interview Data from Survey Pilots and Learning Excerpts
The interview data was analyzed thematically by accumulating interview passages into themes and deriving a subsequent interpretation for assessment development. The analytical process took place in reference to Grounded Theory [19]. This qualitative method combines interview data into bundled theoretical interpretations justified by the subject matter. Interview data was coded and then assigned to categories [11] to gain an overview of respondents’ assessment experience. The following two themes are exemplary as categories used for further survey development, which are based on the coded interview data:
Issues of Question Comprehensibility. In the first pilot phase, all respondents pointed out unclear question phrasing, with some questions rereading. This led to specific rephrasing during the first iterative revision. In the second pilot phase, two respondents mentioned unclear phrasing, and stagnant rereading only occurred in the case of one respondent.
Engagement With Text Fields. Dimension 4 on Critiquing consisted of, among other questions, two optional text fields per visualization example. The survey pilots showed varying attitudes towards expression in text form. Three out of eight respondents principally skipped text fields; two articulated some thought; another two did not fill in the fields, even though they verbally expressed suitable inputs to the present researcher. The eighth respondent filled out all text fields conscientiously to, as was self-stated, perform well in the survey. So that no assumptions are made about the respondent’s expressive ability or their response attitude, text inputs as indicators for critical ability were kept optional. Later in this paper, the text inputs (Q4d/Q4e) are evaluated as additional insight into critical ability (in Section 5.4).
4.3 Quantitative Testing: Representative Survey
After the qualitative survey pilots, the assessment questions and the survey design were iterated to present a validated version of MAVIL to a survey panel. The panelists represented a German-speaking country’s age group distribution from 18 to 74 years and the male-female gender split within those age groups. This survey was implemented with the aid of a market research organization. Valid survey cases had to pass certain quality criteria and the resulting data set, evaluated in Section 5, held 438 cases.
4.3.1 Data Collection in Cooperation With a Market Research Organization
So that MAVIL could undergo an implementation with a representative scope, respondents were recruited in cooperation with a market research organization. The organization guided respondents to the survey in a controlled manner and pre-selected them based on crossing the agreed-upon quotas of age groups and gender. The respondents originated from the organization’s database or other websites promoting market research participation. Respondents received incentives, whose form depended on the panelist’s affiliation with either the market research organization or other platforms. The organization held demographic self-reported data from the panel, the accuracy of which was verified by collecting demographic data in our survey.
During the cooperation with the market research organization, 457 respondents completed the assessment over 14 days via an online survey. The number of completed surveys was achieved in subsets of 50 completions. We evaluated the data quality of the completed surveys after each subset, followed by the criteria for valid survey cases, as illustrated in the next subsection.
4.3.2 Quality Criteria for Valid Survey Cases
The completed survey cases were downloaded as a data set from SoSci, where the survey had been self-administrated on our part. All 457 cases from the market research cooperation period from March 3 until March 17, 2023, were selected. A completed run held responses for all obligatory questions (i.e., excluding the optional text fields), and both tested data visualizations. Respondents who exited the survey early were omitted. Another 19 cases were disregarded, as these respondents failed to answer the attention question (MAVIL Q5AT, see Figure 1) correctly. Referring to the surveyed bar chart (2(b)), this question asked respondents to name the first country listed in the $\text{CO}_{2}$ emissions chart. The correct answer was China, with variations like "china" or "Chin" also accepted.
The resulting data set, consisting of completed surveys during which the attention test was passed, shows 438 valid cases. The survey data set will be made openly accessible upon publication of this paper.
4.3.3 Demographics of the Survey
Demographic information on the respondents included their gender, age group, and education level. Respondents also had to confirm their country of residence where the survey was conducted. Inclusive options for gender self-assignment were offered. Of the 438 respondents, 223 were male ($50.9\%$) and 214 ($48.9\%$) female. One respondent reported their gender as inter. Age groups were recorded in categories of 10-year ranges and represented the distribution of the surveyed population. The minimum registered age was 18 since the legal age was a consent criterion, and the maximum was 74 years, as this was the highest age of online panelists available to the market research organization.
The representation of education levels, though meaningful, was not sought after because, among other things, the standardized procedure of the market research organization did not provide for the representation of this demographic feature. The sample turned out to over-represent respondents with higher education, such as university degrees.
5 Results
MAVIL was tested in an online survey with $438$ respondents representative of a German-speaking country (age groups 18 to 74 years and male-female gender split). In the following, the results from this implementation are showcased, as they provide the first indications for the feasibility of the assessment. The survey data is analyzed to share insights into the ability composition of the surveyed population and the understanding of a line and bar chart showing climate data (see Figure 2).
5.1 Insights into Performance Measures
To give a general impression about completing the assessment: It took respondents an average of $2.75$ minutes to answer the main body of MAVIL questions per data visualization. On average, framing elements of the survey, such as the introduction and the query of demographic info, took about a minute to complete.
Based on the data reading tasks posed in MAVIL Dimension 5, there were clear deficits when evaluating the surveyed population’s Visual Data Literacy from a task-based point of view. Overall, $48\%$ of the respondents made mistakes when reading the simple line and bar chart. Effect sizes of differences between the two charts are relatively small, as seen when comparing the answer distribution of correct answers (see Figure 3). The results of the data reading tasks in Dimension 5 across the visualizations were summarized into a Data Reading Score (DRS), which corresponds to the sum of all correct data reading answers per respondent (see Table 2).
Figure 3: Data reading tasks (Q5a-Q5d) were presented in true-or-false mode (see Figure 2). In this chart, the percentages of correct and false responses given are shown, the total number of answers being 438.
Line Chart
2 Correct
Line Chart
1 Correct
Line Chart
0 Correct
Bar Chart
2 Correct
52% 18% 3%
Bar Chart
1 Correct
13% 5% 2%
Bar Chart
0 Correct
4% 3% -
Table 2: Respondents (in %) with correct answers for the data reading questions.
5.2 Respondents’ Self-Assessed Abilities and Understanding of the Surveyed Charts
In contrast to the data reading performance, a considerably smaller portion of the respondents indicate that they are not able to summarize the content of the shown data visualizations (MAVIL Q1c, see Figure 1). In the case of both charts, only $5\%$ of respondents state that they cannot summarize the visualization. This suggests the possibility of misleading information transfer, as many respondents are unaware of reading deficits. Regarding the displayed chart types (Q2c), $20\%$ of respondents did not indicate familiarity with the line chart, and $19\%$ did not indicate familiarity with the bar chart. The intersection of these statements results in $12\%$ of respondents who felt unfamiliar or uncertain with the line and the bar chart as graph types. Further, respondents have self-assessed deficits in understanding simple measurement units (Q2a). $22\%$ of respondents did not fully comprehend the usage of millimeters, and $28\%$ of respondents found millions of tons somewhat incomprehensible.
There was a notable difference in the aesthetic perception of the surveyed climate data visualizations. Around $10\%$ of respondents rated the impression of the bar chart more positively than that of the line chart. In particular, $60\%$ agreed to find the surveyed bar chart pretty (Q3b), and $72\%$ found it well-designed (Q3c). The line chart was perceived as slightly less aesthetic, with $58\%$ agreeing to find it pretty and $68\%$ finding it well-designed. This information tends to agree with the assessment of visualization trustworthiness (Q3d), as $75\%$ of respondents found the bar chart trustworthy, with a negative five percent difference from those who found the line chart trustworthy.
5.3 Correlative Exploration of MAVIL Variables
The correlative exploration of MAVIL dimensions with scale ratings supports the creation of distinct ability dimensions: MAVIL questions within Dimension 1 through Dimension 4 show steady inner cohesion (see 4(a)) and simultaneously, these dimensions differ from one another, suggesting they query distinctive ability areas (see 4(b)). The correlative exploration highlights some variables, such as self-assessed numeracy, as significant indicators for Visual Data Literacy.
Statistical analyses were conducted using Python (with statistical tests from the scipy.stats) module and R (using the MASS and psych packages). Alongside non-parametric tests, we also used parametric methods, following arguments that they are appropriate for Likert-scale data even with small sample sizes, unequal variances, or non-normal distributions [64]. The correlations (in Section 5.3.1) were calculated according to Pearson’s correlation coefficient using the median whenever results had to be aggregated over dimensions or visualizations. To cross-check and thus confirm our statistical results, we also calculated Spearman’s rank correlation coefficient for the questions within each dimension and found that the results were almost identical with deviations between the two correlation coefficients below $0.05$. For the binary results (in Section 5.3.2), a t-test for the arithmetic mean and Mood’s median test were conducted and gave similar results. Both tests gave the same result with regard to rejecting the null hypotheses. Additionally, we looked at ordinal regressions, with the outcomes supporting the hypotheses (see Figure 5).
(a) Between questions (Dim. 1–4) by visualization example.
(b) Between dimensions (Dim. 1–4) and DRS.
Figure 4: Pearson correlation tables for MAVIL. (a) shows correlations between individual questions (Dimensions 1–4), split by the visualization example shown. (b) displays correlations between MAVIL ability dimensions and DRS (based on Dimension 5), aggregated across both examples. All correlations are statistically significant ($p<0.05$). See supplementary Figures S1–S3 for detailed item-level views.
5.3.1 Correlating Assessment Variables from MAVIL
This part of the correlative analysis explores relationships between MAVIL dimensions as well as between individual questions within each dimension, using statistical methods and significance testing. A consistent positive correlation of questions within Dimension 1 through Dimension 4 indicates the plausibility of constructing ability dimensions. This holds true for both Pearson and Spearman correlation coefficients. As shown in the correlation tables in 4(a) (based on Pearson), the relations between questions within dimensions for both surveyed data visualizations are fairly strong. For Dimension 1 (Comprehending) and Dimension 2 (Decoding), all coefficients fall within the moderate (0.3-0.49) to strong (0.5-0.69) range. For Dimension 3 (Aestheticizing) and Dimension 4 (Critiquing), correlations are strong to very strong (0.7-1). Note that, for Dimension 4, only the questions of Likert type (Q4a-Q4c) were considered for the correlation and that Q4d and Q4e, which were optional text fields, are evaluated separately (see Section 5.4).
To evaluate internal consistency, we also calculated Cronbach’s alpha for Dimensions 1-4, which ranged from $0.75$ to $0.9$, indicating good to excellent reliability. These results further support the aggregation of items within each dimension. Given the distributional properties of the data, we opted to use the median as a robust summary measure.
When dimensions were fully grouped across questions and visualizations, as done in 4(b), Dimension 1 on Comprehending strongly correlates with Dimension 2 on Decoding and Dimension 3 on Aestheticizing. Their correlation with the Data Reading Score (DRS), which corresponds to the sum of correctly solved data reading tasks per respondent, varies and remains weak (0.1-0.29) to moderate. Dimension 4 on Critiquing disrupts this dynamic, as it is the only dimension correlating negatively with Dimension 1 through Dimension 3 and the DRS. This could be due to the simplicity of the data visualizations, as they might limit the scope of finding errors (Q4a) or coming up with design suggestions (Q4c). Respondents might react differently if hidden errors or intentional design issues were to be part of the assessed visualizations.
5.3.2 Significance of Age and Numeracy
Age and numeracy correlated significantly with the number of correctly solved data reading tasks. Binary t-tests and Mood’s median test on MAVIL variables, including demographic features, were carried out in relation to the Data Reading Score (DRS). They showed that the youngest age group of the surveyed population outperformed the older respondents regarding their DRS. The age group from 18 to 24 years ($14\%$ of respondents) had a DRS median of 4 (with p=0.01). The age groups from 25 to 74 years ($86\%$ of respondents) had a DRS median equal to 3 (with p=0.01). In addition, the self-assessed numeracy level (Q6b) correlates significantly with the DRS. Those respondents who rated their numeracy level high ($37\%$ of respondents) tended to have a DRS median equal to 4 (with p=0.009). Those who rated their numeracy level neutral or low ($63\%$ of respondents) had a lower DRS median equal to 3 (with p=0.009), suggesting clear value of having respondents self-assess their ability level.
We further validated these results using ordinal regression. The outcomes showed a substantially higher probability of achieving a DRS of 4 out of 4 in the youngest age group compared to other age groups (see Figure 5(a)). Additionally, the regression confirmed the influence of self-assessed numeracy on data reading performance, as shown in the predicted probabilities in Figure 5(b).
There is no correlation between the highest level of education reported and the self-assessed level of numeracy (Pearson’s r=0.016). As expected, our analysis did not observe a meaningful difference in Visual Data Literacy dimensions based on gender.
(a) Probability for each DRS depending on the age category of the respondents. Younger respondents had a higher probability of answering all data reading tasks correctly.
(b) Probability for each DRS depending on the numeracy self-assessment of the respondents. Higher self-rated numeracy increased the likelihood of correctly answering all data reading tasks.
Figure 5: Ordinal regression supports the observed relationship between age, self-assessed numeracy, and data reading task performance.
5.4 Report on Visualization Criticism in the Text Inputs
For each visualization, respondents had the option to provide written visualization criticism in two text fields (max. $2,000$ characters each), one for visualization strengths (Q4d) and one for visualization weaknesses (Q4e). These questions were not mandatory, yet $37\%$ submitted at least one meaningful response. All inputs for the line chart and the bar chart were categorized into i. Visualization Strengths, ii. Visualization Weaknesses and iii. Climate Commentary.
5.4.1 Categories of Text Inputs and Examples
i. Visualization Strengths were queried in Question Q4d and defined in our analysis as a category of text inputs where a clear positive point of critique related to the tested visualization was made. For the Line Chart (see 2(a)) $99$ inputs were given, such as finding the chart "inviting, easy to understand, the color blue well chosen", or that noting it was "kept quite simple not cluttered". Someone else highlighted that "the source is indicated, [which] must be highly credited". For the bar chart (see 2(b)) $113$ respective inputs were provided, describing it as "easy to understand and memorize", acknowledging that it "raises awareness", or finding "described data easily comparable".
ii. Visualization Weaknesses could be articulated for Q4e and in our analysis were defined as comments highlighting clear negative points of critique related to the tested visualization. For the line chart, $87$ inputs were provided, stating, for example, that "the graphic looks outdated, could be designed better". A few respondents were "missing axis labeling" or wondered what the uncertainty margin means. For the bar chart, there were $78$ inputs, including debate on the cloud icon, finding it "unnecessary" or "strange" (though this graph element had also been highlighted as a strength, e.g., someone said that the "cloud speaks to [them]"). Others criticized the comparison of countries, and, for instance, found the "relation to population missing" or wanted more information because "there are other countries worth mentioning that also emit high levels of CO2".
iii. Climate Commentary appeared in Q4d and Q4e and emerged in our analysis as opinions about and reactions to climate change, rather than comments on the tested visualization itself. In total, there were $35$ of such inputs. To give an impression of this category, one respondent wrote that "the so-called man-made climate change is pure fraud", others called climate change "a lie" and one respondent said that the visualization would aim towards "tricking people into fear". Others articulated feeling affected such as that the visualization "makes [them] sad" or that looking at it they felt that "the world will be destroyed".
5.4.2 Statistical Analysis of Text Inputs
A chi-squared test was employed to evaluate the relationship between the non-ordinal text inputs with demographic factors and respondents’ Data Reading Scores (DRS). For this test, i. Visualization Strengths ($212$ inputs total) and ii. Visualization Weaknesses ($165$ inputs total) were considered equally by coding them as $1$s and putting them in a 2x2 contingency table for each data visualization. Due to partially small sample sizes, significant findings from the chi-squared test were validated through an additional Monte Carlo simulation [24, 74] by randomly sampling $2,000$ replicates.
For both the line and bar chart, results indicated a significant difference in text input provision across education levels (Line Chart: ${X}^{2}$=12.93, df=1, p=0.0003, Bar Chart: ${X}^{2}$=7.88, df=1, p=0.005). The simulated p-values from the Monte Carlo simulation, based on $2,000$ replicates, validate the significance of this finding (Line Chart: p=0.0005, Bar Chart: p=0.005). This suggests that respondents with secondary and tertiary educational degrees were the ones who tended to provide optional visualization criticism. As educational disparities can impact survey responses [22], possibly, higher-educated respondents may be more accustomed to survey items demanding written input and have a structural advantage in providing this type of survey answer.
Moreover, a significant association was found between the provided text inputs and the DRS (Line Chart: ${X}^{2}$=10.40, df=1, p=0.001, Bar Chart: ${X}^{2}$=8.91, df=1, p=0.003). This association was validated through a Monte Carlo simulation with $2,000$ replicates (Line Chart: p=0.001, Bar Chart=0.002). These findings indicate that the respondents with a higher DRS of 3 or 4 were the ones who were more likely to articulate visualization criticism. This underscores the position that high critical ability is linked to a higher level of literacy [17]. In the chi-squared tests, no significant differences were found regarding the association of text input provision across neither age groups nor gender.
For the subset of 20 respondents who provided inputs categorized as Climate Commentary (35 inputs in total), we applied a binary coding scheme (Climate Commentary: yes/ no) and conducted an independent samples t-test on the overall dataset. The results showed a mean DRS of $3.32$ for those without climate commentary and $2.9$ for those who shared an opinion, however, they were not statistically significant ($p=0.097$). While the full-sample t-test results are almost statistically significant, this appears to be driven by the imbalance in group sizes. When accounting for this imbalance through a bootstrap-like approach, the effect was not robust ($p\approx 0.23$), suggesting that the observed difference may be sensitive to sample composition rather than reflecting a consistent underlying difference.
Given the small and uneven sample sizes, respondents’ attitudes toward climate change may warrant closer examination in future research. These inputs highlight the salience of the climate discourse and that they may be valuable when analyzing respondents’ engagement with visual information, as emotional states can influence how information is extracted [46].
6 Discussion
With MAVIL, we propose a broader lens for assessing Visual Data Literacy—one that acknowledges both performance, self-perceived ability ratings, and written critique. While our goal is not to replace assessment tools like VLAT [78] or CALVI [34], MAVIL challenges established assessments by introducing additional ability dimensions and integrating underutilized measurement approaches, such as self-assessed ability ratings. This multidimensional approach may offer insights into how general audiences not only read but also perceive and reflect on their engagement and abilities related to data visualizations—insights that task-based methods alone cannot fully capture.
In this section, we reflect on the conceptual and practical limitations of our case-specific implementations of MAVIL, discuss what our survey results reveal about different dimensions of literacy, and propose directions for expanding MAVIL’s applicability. Although we focused on two simple, static climate visualizations, the results offer a compelling entry point to better understand how audiences engage with real-world data visualizations and what assumptions researchers, designers, or anyone who creates or shares visualizations might need to rethink.
6.1 Survey-Based Insights from MAVIL
Traditional assessments [78, 34] excel in task-based performance, such as value retrieval or pattern identification, but overlook other dimensions of how audiences perceive and interpret visualizations. We gathered representative survey data for a German-speaking country (N=$438$) and collected insights specific to two climate data visualizations across six newly constructed Visual Data Literacy dimensions. We assessed not only performance through data reading (MAVIL Dimension 5) but also perceived comprehension (Dim1) and decoding ability (Dim2), aesthetic judgment (Dim3), critique behavior (Dim4), and context-related self-assessments like numeracy and topic interest (Dim6). While findings are specific to the visualizations used, they offer valuable implications for visualization design, information communication, and the understanding of critical data by general audiences. Even simple, static charts still pose challenges, which is exemplified by $12\%$ of respondents feeling unfamiliar with both line and bar charts (see Section 5.2).
The MAVIL survey further revealed that roughly a quarter of respondents felt unable to decode data units like millimeters and millions of tons, highlighting that elements implicitly assumed to be self-explanatory by designers or curators [48] may require reconsideration. This highlights that correct interpretation by general audiences cannot be taken for granted, underscoring the value of identifying friction points in audience understanding, which can in turn inform visualization design and implementation [47]. Such findings also call into question one-size-fits-all approaches, which have been criticized for being ineffective with diverse users who have different literacy levels and types of knowledge [49, 75].
Including aesthetic perception in our literacy assessment revealed that respondents who found a visualization visually appealing also tended to trust it (see Section 5.2), suggesting that, in line with existent literature [23], design and perceived credibility are closely linked. This underscores the importance of further investigating which visual elements or design choices different audience groups may perceive as appealing or trustworthy. For example, proximity techniques have been shown to increase viewer engagement [14, 46], and such insights could guide the creation of visualizations that are both inviting and credible. However, in our study, aesthetic perception had limited influence on actual interpretation, as this dimension showed only a weak correlation with data reading performance (see 4(b)).
Respondents’ open-ended visualization critiques (see Section 5.4) offered valuable qualitative insight and illustrated their capacity for critical engagement. In Dimension 4, respondents provided a total of $377$ points of critique on visualizations’ strengths and weaknesses, with some comments highlighting distracting design choices or missing labels, offering grounded insight into what viewers notice and what may hinder understanding. However, in their self-ability ratings most respondents did not feel confident suggesting design improvements, indicating a gap between recognizing and articulating critique. Likely, more flawed or intentionally misleading visuals (for instance done by Camba et al. [20] or Ge et al. [34]) would elicit stronger rating responses, and future assessment implementations could explore this space.
6.2 Limitations
As with any study, certain limitations should be acknowledged. These include aspects related to the surveyed population, such as representation boundaries. Respondents below the age of $18$ and above the age of $74$ were not considered in the survey because of legal age as a consent criterion and due to availability reasons (see Section 4.3.3). The respondents share the commonality of being incentivized by a market research organization, which may introduce a bias toward individuals with frequent online survey participation and, therefore, higher digital skills. However, the approach we have chosen to select respondents is also standard in other representative surveys (e.g., [15, 39, 16]).
We recognize that assessment results are likely to vary based on respondents’ socio-cultural and geographical backgrounds, which limits the scalability and comparability of MAVIL. While it allows for inputs from different perspectives, including optional open-ended responses, MAVIL has not been tested for transferability across other cultural or contextual settings. Instead, we used MAVIL to support context-specific interpretation and demonstrated in our representative survey (see Section 4.3) that age and self-perceived numeracy significantly influenced data reading performance (see Section 5.3.2).
MAVIL’s integration of self-perceived ability ratings offers valuable insights into how people engage with visualizations beyond task-based performance, but it may potentially introduce sources of noise. Self-assessments can be influenced by factors like over- or underestimation, social desirability, or ambiguity in question interpretation [36, 57]. While we have piloted MAVIL in several qualitative rounds to ensure clarity and ease of interpretation (see Section 4.2), biases inherent to self-assessment may remain. Nevertheless, as discussed in the Background (see Section 2.3), self-perceived ability ratings are a recognized approach in the study of abilities and continue to provide valuable insights [77, 36, 50].
The visualizations used in the survey, a line chart and a bar chart on climate data, were chosen for their visual simplicity and assumed broad familiarity (see Section 4.3.2). This helped ensure accessibility to general audiences but may limit the scope of our insights. Two visualizations cannot be representative of the variety, complexity levels, and design possibilities of data visualizations. MAVIL should be tested with more diverse visualizations, such as interactive visualizations or intentionally misleading ones, to evaluate how it performs under varying complexity and design conditions.
We deliberately selected climate visualizations for their real-world relevance, aiming to assess how people engage with meaningful, high-stakes information rather than artificial stimuli (see Section 4.1.1). While such topics may evoke emotionally charged or politically influenced reactions, we argue that these responses are a valid part of the interpretation process and reflect how visualizations are encountered in practice. Studying potentially polarizing visuals is therefore, not only justified but essential for understanding real-world information retrieval and addressing misinformation [55, 26]. To account for possible bias, we let respondents assess their topic interest (see MAVIL Q6a in Figure 1). We also examined unsolicited climate-related commentary in our qualitative data and tested whether these participants differed in their data reading performance, finding no significant effect (see Section 5.4.2). That said, we acknowledge that topic-related influences, whether related to climate or other subjects, would benefit from further investigation. This could mean including additional questions on topic interest to provide more nuanced contextualization of responses in future surveys.
6.3 Future Work
Future implementations of MAVIL should be expanded to include a wider variety of visualization formats. Beyond static charts, interactive visualizations are increasingly common in digital communication, possibly requiring different skills [45] and forms of audience engagement [29]. Testing MAVIL on such formats may warrant an additional dimension to assess interaction-based ability facets. Other common visualization types, including infographics, could be explored to reflect broad visual data communication practices [51, 44, 37, 38].
MAVIL could further benefit from integration with existing tools like Mini-VLAT [67], as a combined assessment would allow for further comparison of self-assessed and performance-based results within the same sample. Initial correlations we observed, such as between self-assessed numeracy and data reading performance, highlight promising grounds for such work. While this study was focused on climate data, MAVIL could be adapted to other domains by rephrasing topic specificity in the Contextualizing dimension, offering a promising direction for future work on assessing its transferability.
Finally, MAVIL could incorporate an educative feedback component. Respondents frequently expressed curiosity about their performance in the data reading tasks during qualitative testing. Providing immediate feedback, correct answers, or even a summary of self-assessed literacy strengths and weaknesses could enhance the assessment’s value, encourage reflection, while serving literacy-building [17], and could function as an incentive for respondents.
7 Conclusion
With MAVIL we contribute a novel approach to assessing Visual Data Literacy by conceptualizing it as a multidimensional ability composed of distinct facets that can be partially self-assessed. Grounded in a theory-driven approach, it introduces six ability dimensions and combines task-based, self-assessed, and open-ended question formats. This expands existing literacy assessments by capturing not only performance measures but also self-perceived ability ratings. We tested MAVIL in what is, to our knowledge, the first representative survey to assess Visual Data Literacy using a multidimensional approach. Validation was supported through a nationally representative sample covering age groups $18$ to $74$ and a balanced male–female gender split. Our MAVIL survey provided nuanced insights into respondents’ Visual Data Literacy and their understanding of two climate data visualizations. The findings revealed self-assessed deficits, including perceived challenges in interpreting simple measurement units and some unfamiliarity with presumably simple chart types. Overall, MAVIL assesses a combination of different yet central aspects of visualization understanding and aims to encourage more nuanced and critical approaches to evaluating Visual Data Literacy in general audiences.
Figure Credits
  • 2(a): Statista, Der Meeresspiegel steigt kontinuierlich an, NASA, https://de.statista.com/infografik/21922/anstieg-des-meeresspiegels/
  • 2(b): Statista, Diese Länder stoßen am meisten $CO_{2}$ aus, Global Carbon Project, https://de.statista.com/infografik/18287/co2-emissionen-in-ausgewaehlten-laendern/